OpenStudy (anonymous):

Find the value of y that makes the statement true. 14 – 3y = 4y A. 12 B. 2 C. –12 D. –2

OpenStudy (unklerhaukus):

to solve 14–3y=4y for y add 3y to both sides of the equation then divide both sides of the equation by 7

OpenStudy (anonymous):

thank u

OpenStudy (unklerhaukus):

what did you get?

OpenStudy (anonymous):

2

OpenStudy (unklerhaukus):

\[\Large\color{red}\checkmark\]

OpenStudy (unklerhaukus):

if you can to check your solution you can do it this way 14 – 3(2) = 4(2) 14 – 6 = 8 8 = 8 which is a true statement so y=2 IS a solution to the original equation